OpenStudy (anonymous):

if h(x)=sqrt(2+2f(x)) where f(1)=7 and f'(1)=2 find h'(1)

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@aaronq @Directrix

OpenStudy (phi):

can you find the derivative of \[ \sqrt{2+2f(x)} \] ?

OpenStudy (anonymous):

no

OpenStudy (phi):

do you know the power rule for taking the derivative of \( x^n\) ? \[ d x^n = n x^{n-1} dx \] re-write the square root as \[ (2+2f(x))^{\frac{1}{2}} \] and use the power rule

OpenStudy (anonymous):

yeah the power rule is nx^(n-1)

OpenStudy (phi):

the power rule works even if n is a fraction. anything to the 1/2 power is the same as the square root

OpenStudy (anonymous):

ok so it would be (sqrt(2+2f(x)))/2

OpenStudy (phi):

\[ (2+2f(x))^{\frac{1}{2}} \] the rule says "bring down" the 1/2 so you will get times 1/2 then make a new exponent = 1/2 - 1 you did not do that part

OpenStudy (anonymous):

ok it would be -(sqrt(2+2f(x)))/2

OpenStudy (phi):

first, get rid of the square root. use to the 1/2 power instead \[ (2+2f(x))^{\frac{1}{2}} \] now apply the rule \[ d x^n = n\ x^{n-1} dx \] in your case, n = 1/2 and x is (2 + 2 f(x) )

OpenStudy (anonymous):

ok i am lost. what would it be then

OpenStudy (phi):

Are you asking how to interpret the power rule ? \[ d x^n = n\ x^{n-1} dx \] says, to take the derivative of x^n write n out front. then find n-1, and write \( x^{n-1}\) then take the derivative of x try those steps on \[ (2+2f(x))^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[(1/2)(2+2f(x))^{-1/2}\]

OpenStudy (phi):

yes, but the last step is take the derivative of x, or in this case \[ d ( 2 + 2 f(x) ) \]

OpenStudy (phi):

d(2 + 2 f(x)) = d(2) + d (2 f(x) ) derivative of a constant 2 is 0. d (2 f(x)) is 2 * d (f(x)) we don't know the details of f(x) so that is the answer: d (f(x)) d( f(x) ) is also written f'(x) (it means the same thing)

OpenStudy (phi):

so you now have \[ \frac{1}{2}(2+2f(x))^{-1/2} 2 f'(x) \] which you can simplify to \[ (2+2f(x))^{-1/2} f'(x) \] or if you don't like negative exponents, you can "flip" it and wriate \[ \frac{f'(x)}{(2+2f(x))^{\frac{1}{2}}}\] or if you would rather see the square root instead of the 1/2 power \[ \frac{f'(x)}{\sqrt{2+2f(x)}} \]

OpenStudy (phi):

now you can find the answer using f(1)=7 and f'(1)=2 find h'(1)

OpenStudy (anonymous):

so its 2/sqrt30

OpenStudy (phi):

you have \[ h'(x) = \frac{f'(x)}{\sqrt{2+2f(x)}} \] you know f(1) =7 and f'(1)=2 so you can find h'(1)

OpenStudy (anonymous):

i got 2/sqrt30

OpenStudy (phi):

what is 2+ 2*7

OpenStudy (anonymous):

never mind its 2/sqrt16=1/2

OpenStudy (phi):

much better.